The perimeter of a rectangle is 76 m. If the width were doubled and the length were increased by 5m, the perimeter would be 120 m. What are the length and width of the rectangle?
Answer by ankor@dixie-net.com(22740)

The perimeter of a rectangle is 76 m.
If the width were doubled and the length were increased by 5m, the perimeter would be 120 m.
What are the length and width of the rectangle?
:
2L + 2W = 76
Simplify divide by 2
L + W = 38
or
L = 38 - W
:
"If the width were doubled and the length were increased by 5m, the perimeter would be 120 m."
2(L+5) + 2(2W) = 120
2L + 10 + 4W = 120
2L + 4W = 120 - 10
2L + 4W = 110
Simplify divide by 2, subtract the 1st equation
L + 2W = 55
L + W = 38
-------------Subtraction eliminates L, find W
W = 17 is the width
then
L = 38 - 17
L = 21 is the length
;
:
Check solution in original perimeter equation
2(21) + 2(17) = 76